OpenStudy (cderaalien):

I was trying to the variable 'y' isolated in the following expression, x + y = 2. I did the following: I subtracted 'x' from both sides, and I came up with y = -1x -2. Did I do it right? (I was trying to isolate the variable 'y'.)

OpenStudy (cderaalien):

ah my favorite, @pooja195

pooja195 (pooja195):

xD ily2 <3 Yes that's correct

pooja195 (pooja195):

You can also write it as y=-x+2

OpenStudy (cderaalien):

I know, I just get confused sometimes when it's just 'x' and I forget the meaning of just the 'x' variable, even though in the back of my mind, I know it's just 1.

OpenStudy (cderaalien):

Thank you. <333

pooja195 (pooja195):

Yep that's what i do but good job well done ^-^ <3

pooja195 (pooja195):

yw ^_^

OpenStudy (cderaalien):

Can I ask something else?

pooja195 (pooja195):

go ahead :-)

OpenStudy (cderaalien):

How did you become so good at math, as in remembering the material? I have the ISEE test coming up and I'm super nervous, because it contains geometry and algebra 1, etc.

pooja195 (pooja195):

Practice, helping users on this site keeps my memory fresh I can help with anything up to Geometry since that's as far as i have gotten.

OpenStudy (cderaalien):

Also, why did you put the addition sign, when it's a minus sign (from subtracting x from both sides)

pooja195 (pooja195):

the 2 is positive

pooja195 (pooja195):

im not sure what ur referring to :-/

OpenStudy (cderaalien):

Okay, so I asked if y = -1x - 2 was the right answer. You said, "yes, it can also be written as "y = -x +2"

OpenStudy (cderaalien):

Why is there a "+2" instead of a "-2"?

pooja195 (pooja195):

\[\huge~\rm~ x + y = 2\] We are isolating only y \[\huge~\rm~ x-x + y = 2-x\] By subtracting x from both sides we get the equation \[\huge~\rm~\bf~y=2-x\] We can leave it like this or we can write it like this \[\huge~\rm~y=-x+2\] To begin with the 2 was always positive

OpenStudy (cderaalien):

ah, so just because someone is subtracting, it wouldn't change the positive 2 to a negative 2?

pooja195 (pooja195):

Correct 2 is not a like term now if it was 2x we would subtract an x but 2 is not a like term so we dont subtract from it

OpenStudy (cderaalien):

Thank you so much, sorry for not understand!! You're one of the best! Thank you again!!

OpenStudy (cderaalien):

not understanding *

pooja195 (pooja195):

thank you and yw ^_^ It's alright that you don't understand i'm happy you are asking questions! It shows that you are willing to learn and that's all I care about :-)